# DynFOA: Generating First-Order Ambisonics with Conditional Diffusion for Dynamic and Acoustically Complex 360-Degree Videos

### v1

@@ Abstract
Introduction Spatial audio is a cornerstone of immersion in Virtual Reality (VR) and 360-degree media, allowing users to perceive sound with realistic direction, distance, and environmental context [spatial_audio2023]. Unlike conventional stereo audio, spatial audio aligns the auditory environment with visual cues, enabling a coherent perceptual scene. However, generating high-fidelity, physically plausible spatial audio remains challenging. Realistic spatial audio requires not only precise synchronization with visual events but also accurate 3D localization of sound sources, realistic modeling of occlusion, reflections, and reverberation, and adaptation to the listener’s head orientation in real time [vr_audio2022].
A major limitation in existing works is the under-utilization of the rich geometric and semantic information available in 360-degree video for audio rendering. Many methods, such as some FOA-based approaches or hybrid audio-visual models, focus primarily on sound source localization without modeling how the surrounding environment shapes the sound field. For instance, OmniAudio [liu2025omniaudio] generates FOA audio but assumes static, context-free sound sources, while Sonic4D [sonic4d2025] estimates 3D positions but outputs only stereo audio, reducing immersion. These approaches typically neglect how environmental factors, such as moving objects, walls, or furniture, affect propagation through occlusion, reflections, and reverberation. This omission results in sound fields that lack physical grounding and fail to adapt convincingly to user orientation.
@@ 1.1 Spatial Audio Representations and Rendering
Spatial audio reproduction in VR has long relied on Ambisonics, which encodes sound fields in the spherical harmonic domain [gerzon1973periphony]. First-Order Ambisonics (FOA) strikes a balance between spatial resolution and computational efficiency, while higher-order Ambisonics improves angular precision at a higher computational cost. The mathematical foundation of spherical harmonic decomposition allows for 3D sound field capture and reconstruction at varying levels of accuracy, making Ambisonics particularly well-suited for VR, where perceptual fidelity must be balanced with real-time processing constraints.
For spatial audio rendering, Head-Related Transfer Functions (HRTFs) are essential, providing the necessary filtering to simulate spatial localization. Standardized resources, such as KEMAR [gardner1994kemar] and CIPIC [algazi2001cipic], along with exchange formats like SOFA [sofa2022], form the foundation for HRTF-based rendering. However, individualization remains a challenge due to anthropometric variability, which affects localization and externalization of sounds. Recent works have addressed this issue through both direct measurements and computational models. Platforms for acoustic simulation, such as SoundSpaces [chen2020soundspaces, chen2022soundspaces2], offer 3D scanned environments with physically-grounded propagation models, enabling systematic evaluation and cross-modal learning. These platforms emphasize the importance of jointly modeling visual and acoustic cues for better audio-visual scene understanding. In the context of 360-degree video, early work [morgado2018selfsupervised] generated FOA from mono audio, using panoramic cues. OmniAudio [liu2025omniaudio] formalized the 360V2SA task, but with the assumption of static sources and negligible occlusion or reverberation. Sonic4D [sonic4d2025] reconstructed dynamic 3D trajectories for viewpoint-adaptive binaural rendering, though it did not produce FOA. More recent geometry-aware methods [parida2022wacv, points2sound2022] leverage depth cues, but they typically focus only on source distance.
@@ 1.2 Cross-Modal Learning for Sound Localization and Separation
Cross-modal learning leverages the synergy between different modalities: visual signals help constrain sound localization and disambiguate overlapping events, while audio assists in grounding visual understanding. This synergy is particularly crucial for immersive spatial audio, where multiple sound sources and complex environments necessitate robust sound separation. A significant body of work has explored cross-modal learning for sound localization and separation using visual cues. Weakly supervised methods align objects with sounds [arandjelovic2018objects, senocak2018localize, ramaswamy2020seethesound], while co-separation frameworks learn to disentangle multiple overlapping sound sources. Other approaches focus specifically on speech, such as speaker-independent audiovisual separation [ephrat2018looking] and lip-synchronized speech extraction [gao2021visualvoice]. Large-scale datasets like AudioSet [audioset2017], VGGSound [vggsound2020], and MUSIC [zhao2018soundofpixels] have significantly accelerated these advances by providing diverse audio-visual training resources.
@@ 1.3 Visually Guided Spatialization for 360-Degree Video
Recent advancements have focused on vision-guided audio spatialization. Early approaches addressed mono-to-binaural conversion using neural networks [gao2019mono2binaural], while later works shifted toward stereo or pseudo-binaural generation [xu2021pseudobinaural]. Methods like Points2Sound incorporated geometry and motion cues to improve spatialization [points2sound2022], and newer pipelines introduced explicit depth estimation or scene reconstruction for greater accuracy. Despite these developments, many approaches still oversimplify acoustic propagation, often assuming free-field conditions or modeling only distance attenuation. To address this, OmniAudio [liu2025omniaudio] and Sonic4D [sonic4d2025] made significant progress by targeting panoramic or dynamic sources. However, they did not account for FOA outputs or more complex propagation effects, such as reverberation. Head-tracked playback pipelines further demonstrated the necessity of synchronizing audio rotation with the listener’s orientation, but they remain limited in handling occlusion and material diversity.
@@ 1.4 Diffusion Models for High-Fidelity Audio Generation
Generative models have recently revolutionized audio synthesis. Diffusion models, such as WaveGrad [chen2020wavegrad], DiffWave [kong2020diffwave], and AudioGen [kreuk2022audiogen], have set new benchmarks in waveform generation. Spatial audio frameworks based on diffusion, such as Diff-SAGe [kushwaha2024diffsage] and ImmerseDiffusion [heydari2024immersediffusion], demonstrate strong potential, though they focus primarily on perceptual fidelity rather than ensuring physical scene consistency. Meanwhile, autoregressive models like SoundStorm [borsos2023soundstorm] and controllable music generators [copet2023musicgen] offer sequence-level control but struggle with long-term coherence. Video-to-audio pipelines, such as ViSAGe [kim2025visage] and MMAudio [cheng2025mmaudio], integrate motion and semantic cues but rarely address FOA generation or physically grounded propagation.
@@ 2.1 Problem Definition
The objective of this work is to enable physically grounded and perceptually coherent 4D immersive experiences from 360-degree videos by learning to generate scene-aware spatial audio. Prior approaches to spatial audio rendering often rely on simplified acoustic assumptions, neglecting critical aspects such as dynamic sound sources, concurrent source interactions, and propagation effects including occlusion, reflections, and reverberation. Our framework, DynFOA, directly addresses these challenges by learning from multimodal cues—visual appearance, 3D geometry, and material properties—to synthesize first-order ambisonics (FOA) that faithfully reflect the physical structure and acoustic conditions of the scene [liu2024audioldm2taslp, kreuk2022audiogen].
@@
Solving this problem requires addressing a sequence of coupled sub-tasks across both the visual and audio domains. On the visual side, the model must (i) detect and localize sound-emitting and non-emitting objects, (ii) estimate depth, (iii) perform semantic segmentation, and (iv) reconstruct a geometry- and material-aware 3D representation using techniques such as 3D Gaussian Splatting (3DGS) [yu2023deep, zhao2024perception]. On the audio side, the model must (i) extract directional cues from FOA channels, (ii) encode them into a latent representation, and (iii) model complex propagation phenomena including occlusion, reflections, and reverberation [liu2024audioldm2taslp, xu2023sound].
The key challenge lies not only in localizing sound sources but also in handling multiple, dynamic sources within acoustically complex environments [kim2023spatialaudio]. By grounding audio generation in geometry- and material-aware acoustic descriptors, our model captures both static and dynamic elements of the scene. This enables real-time adaptation to source motion, ensuring accurate localization, separation, and a physically consistent audio field that reflects the spatial relationships inherent in 360-degree visual scenes [chen2023virtualaudio].
@@ 2.3 Video Encoder
The Video Encoder extracts spatial and semantic cues from 360-degree video to support realistic sound propagation modeling and synchronized spatial audio rendering [wang2023360audio, zhang2024visualaudio]. It operates in three stages: (1) sound source localization and depth estimation, (2) semantic segmentation and scene reconstruction, and (3) feature extraction and multi-modal fusion.
@@ 2.3.1 Sound Source Localization and Depth Estimation
The encoder first detects and localizes sound-emitting objects in the scene [li2023dynamic]. Each source $i$ is assigned a bounding box $\hat{b}_{i}$ and an activity score $\hat{y}_{i}$, optimized by:
@@
where $(u,v)$ are image coordinates, $D(u,v)$ is depth, and $\theta,\phi$ are the corresponding elevation and azimuth angles. The resulting point cloud serves as the geometric basis for acoustic modeling. To reconstruct the full scene, a hybrid approach combines Truncated Signed Distance Functions (TSDF) for large-scale structures with 3D Gaussian Splatting (3DGS) for fine details [yang2024hybrid], enabling accurate modeling of occlusion, reflections, and reverberation.
@@ 2.3.2 Semantic Segmentation and Scene Reconstruction
The encoder further applies semantic segmentation to classify scene elements (e.g., walls, floors, furniture). Each class is mapped to frequency-dependent acoustic material properties [schmidt2023acoustic, wu2024material], enriching the reconstructed geometry with absorption and reflection parameters. These semantic and geometric cues are integrated into a 3D scene model [chen20243d], enabling simulation of occlusion, reflections, and reverberation based on both structure and material characteristics. This ensures that environmental effects such as sound blocking, scattering, and decay are faithfully captured.
@@ 2.3.3 Feature Extraction and Fusion
Finally, spatial and temporal features are extracted using Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) [sun2023convLSTM]. These video-derived features are fused with audio representations to jointly model scene dynamics and acoustic conditions. The resulting multimodal features support real-time tracking of dynamic sound sources while incorporating propagation effects (occlusion, reflections, and reverberation), thereby enabling perceptually coherent and immersive spatial audio rendering.
@@ 2.4 Audio Encoder
The Audio Encoder processes FOA signals to extract directional, spectral, and temporal cues that align with the geometry- and material-aware representations produced by the Video Encoder. By coupling FOA features with reconstructed scene properties, it enables context-aware spatialization consistent with the physical layout and acoustic characteristics of the environment [li2023foa, zhang2024audiovisual].
@@ 2.4.1 FOA Extraction and Normalization
From the normalized channels, a CNN extracts compact representations of spectral and directional patterns. Convolutions over the time–frequency domain capture harmonic content and inter-channel correlations, while stacked layers aggregate these into higher-level spatial features [sun2023convLSTM]. The resulting FOA embeddings form the basis for geometry-aware mapping and materially consistent spatial audio generation.
@@ 2.4.2 Spatial and Directional Mapping
To incorporate structural priors, we modulate FOA features with geometric distance and material-dependent absorption [li2023foa]. This accounts for sound attenuation and redirection during propagation:
@@
Directional information is further captured by projecting FOA features onto a spherical harmonic basis, yielding a compact representation of spatial energy distributions [wang2024sht]. This transformation reinforces alignment between audio embeddings and reconstructed scene geometry, enabling accurate reasoning about sound propagation across directions.
@@ 2.4.3 Saliency, Reverberation, and Output
To highlight perceptually relevant cues, an attention mechanism modulates FOA features with visual saliency and acoustic context [zhang2024audiovisual]:
@@
Spatial realism is further enriched by augmenting FOA features with late reverberation profiles that capture long-range energy decay and diffusion [wu2024material, zhou2024material]. Estimated from reconstructed geometry and material properties, these profiles complement direct sound and early reflections, yielding acoustically consistent reverberation patterns.
@@ 2.5 Conditional Diffusion Generator
Inspired from recent spatial audio diffusion frameworks such as Diff-SAGe [kushwaha2024diffsage] and ImmerseDiffusion [heydari2024immersediffusion], we employ a conditional diffusion model to synthesize FOA that remain consistent with reconstructed scene geometry and material properties. Operating in the latent FOA domain, the model integrates structural and acoustic cues—including occlusion, reflections, and reverberation—while accounting for dynamic and multiple sound sources. Starting from a noisy latent representation, a U-Net denoiser progressively reconstructs clean FOA signals over $T$ timesteps. Training follows the denoising diffusion probabilistic model (DDPM) objective:
@@ 2.5.1 Conditioning on Geometry and Material Properties
Reconstructed geometry and material attributes provide the foundation for physically grounded synthesis. The 3D mesh encodes structural layout and surface orientation, while material properties specify frequency-dependent absorption coefficients [wu2024material, zhang2024geomaterial]. Embedding these features allows the model to account for attenuation, diffraction, and spatial filtering effects, thereby ensuring that the generated FOA is consistent with the reconstructed scene. Such conditioning is particularly important in scenarios involving dynamic or multiple sound sources, where accurate modeling of energy interaction with the environment is essential.
@@ 2.5.2 Conditioning for Occlusion, Reflections, and Reverberation
To capture realistic propagation effects, the model is further conditioned on occlusion, reflections, and reverberation. Occlusion features are derived from visibility analysis between listener and sources, modulated by material absorption. Early reflections are estimated by tracing geometric paths, providing echo-like cues that enhance spatial depth. Reverberation is represented using frequency-dependent $T_{60}(f)$ curves, which describe late decay characteristics [heydari2024immersediffusion, liu2024condspatial]. Together, these cues enrich the conditioning stream, enabling FOA synthesis that incorporates both direct sound and its environmental response.

### v2

# arXiv:2602.06846v2[cs.SD] 27 Feb 2026

@@ Abstract
Introduction Spatial audio is a cornerstone of immersion in Virtual Reality (VR) and 360-degree media, allowing users to perceive sound with realistic direction, distance, and environmental context [begault20003]. Unlike conventional stereo audio, spatial audio aligns the auditory environment with visual cues, enabling a coherent perceptual scene. However, generating high-fidelity, physically plausible spatial audio remains challenging. Realistic spatial audio requires not only precise synchronization with visual events but also accurate 3D localization of sound sources, realistic modeling of occlusion, reflections, and reverberation, and adaptation to the listener’s head orientation in real time [brungart2002near].
A major limitation in existing works is the under-utilization of the rich geometric and semantic information available in 360-degree video for audio rendering. Many methods, such as some FOA-based approaches or hybrid audio-visual models, focus primarily on sound source localization without modeling how the surrounding environment shapes the sound field. For instance, OmniAudio [liuomniaudio] generates FOA audio but assumes static, context-free sound sources, while Sonic4D [xie2025sonic4d] estimates 3D positions but outputs only stereo audio, reducing immersion. These approaches typically neglect how environmental factors, such as moving objects, walls, or furniture, affect propagation through occlusion, reflections, and reverberation. This omission results in sound fields that lack physical grounding and fail to adapt convincingly to user orientation.
@@ 1.1 Spatial Audio Representations and Rendering
Spatial audio reproduction in VR has long relied on Ambisonics, which encodes sound fields in the spherical harmonic domain [gupta2010three]. First-Order Ambisonics (FOA) strikes a balance between spatial resolution and computational efficiency, while higher-order Ambisonics improves angular precision at a higher computational cost. The mathematical foundation of spherical harmonic decomposition allows for 3D sound field capture and reconstruction at varying levels of accuracy, making Ambisonics particularly well-suited for VR, where perceptual fidelity must be balanced with real-time processing constraints.
For spatial audio rendering, Head-Related Transfer Functions (HRTFs) are essential, providing the necessary filtering to simulate spatial localization. Standardized resources, such as KEMAR [gardner1995hrtf] and CIPIC [algazi2001cipic], along with exchange formats like SOFA [majdak2013spatially], form the foundation for HRTF-based rendering. However, individualization remains a challenge due to anthropometric variability, which affects localization and externalization of sounds. Recent works have addressed this issue through both direct measurements and computational models. Platforms for acoustic simulation, such as SoundSpaces [chen2020soundspaces, chen2022soundspaces], offer 3D scanned environments with physically-grounded propagation models, enabling systematic evaluation and cross-modal learning. These platforms emphasize the importance of jointly modeling visual and acoustic cues for better audio-visual scene understanding. In the context of 360-degree video, early work [morgado2018self] generated FOA from mono audio, using panoramic cues. OmniAudio [liuomniaudio] formalized the 360V2SA task, but with the assumption of static sources and negligible occlusion or reverberation. Sonic4D [xie2025sonic4d] reconstructed dynamic 3D trajectories for viewpoint-adaptive binaural rendering, though it did not produce FOA. More recent geometry-aware methods [cokelek2025spherical, bhosale2024av] leverage depth cues, but they typically focus only on source distance.
@@ 1.2 Cross-Modal Learning for Sound Localization and Separation
Cross-modal learning leverages the synergy between different modalities: visual signals help constrain sound localization and disambiguate overlapping events, while audio assists in grounding visual understanding. This synergy is particularly crucial for immersive spatial audio, where multiple sound sources and complex environments necessitate robust sound separation. A significant body of work has explored cross-modal learning for sound localization and separation using visual cues. Weakly supervised methods align objects with sounds [zhang2018new, senocak2018learning], while co-separation frameworks learn to disentangle multiple overlapping sound sources. Other approaches focus specifically on speech, such as speaker-independent audiovisual separation [ephrat2018looking] and lip-synchronized speech extraction [pan2022selective]. Large-scale datasets like AudioSet [gemmeke2017audio], VGGSound [chen2020vggsound], and MUSIC [panda2013multi] have significantly accelerated these advances by providing diverse audio-visual training resources.
@@ 1.3 Visually Guided Spatialization for 360-Degree Video
Recent advancements have focused on vision-guided audio spatialization. Early approaches addressed mono-to-binaural conversion using neural networks [parida2022beyond], while later works shifted toward stereo or pseudo-binaural generation [xu2021visually]. Methods like Points2Sound incorporated geometry and motion cues to improve spatialization [lluis2022points2sound], and newer pipelines introduced explicit depth estimation or scene reconstruction for greater accuracy. Despite these developments, many approaches still oversimplify acoustic propagation, often assuming free-field conditions or modeling only distance attenuation. To address this, OmniAudio [liuomniaudio] and Sonic4D [xie2025sonic4d] made significant progress by targeting panoramic or dynamic sources. However, they did not account for FOA outputs or more complex propagation effects, such as reverberation. Head-tracked playback pipelines further demonstrated the necessity of synchronizing audio rotation with the listener’s orientation, but they remain limited in handling occlusion and material diversity.
@@ 1.4 Diffusion Models for High-Fidelity Audio Generation
Generative models have recently revolutionized audio synthesis. Diffusion models, such as WaveGrad [chenwavegrad], DiffWave [kongdiffwave], and AudioGen [kreukaudiogen], have set new benchmarks in waveform generation. Spatial audio frameworks based on diffusion, such as Diff-SAGe [kushwaha2025diff] and ImmerseDiffusion [heydari2025immersediffusion], demonstrate strong potential, though they focus primarily on perceptual fidelity rather than ensuring physical scene consistency. Meanwhile, autoregressive models like SoundStorm [borsossoundstorm] and controllable music generators [copet2023simple] offer sequence-level control but struggle with long-term coherence. Video-to-audio pipelines, such as ViSAGe [kim2025visage] and MMAudio [cheng2025mmaudio], integrate motion and semantic cues but rarely address FOA generation or physically grounded propagation.
@@ 2.1 Problem Definition
The objective of this work is to enable physically grounded and perceptually coherent 4D immersive experiences from 360-degree videos by learning to generate scene-aware spatial audio. Prior approaches to spatial audio rendering often rely on simplified acoustic assumptions, neglecting critical aspects such as dynamic sound sources, concurrent source interactions, and propagation effects including occlusion, reflections, and reverberation. Our framework, DynFOA, directly addresses these challenges by learning from multimodal cues—visual appearance, 3D geometry, and material properties—to synthesize first-order ambisonics (FOA) that faithfully reflect the physical structure and acoustic conditions of the scene [liu2024audioldm, kreukaudiogen].
@@
Solving this problem requires addressing a sequence of coupled sub-tasks across both the visual and audio domains. On the visual side, the model must (i) detect and localize sound-emitting and non-emitting objects, (ii) estimate depth, (iii) perform semantic segmentation, and (iv) reconstruct a geometry- and material-aware 3D representation using techniques such as 3D Gaussian Splatting (3DGS) [samavati2023deep]. On the audio side, the model must (i) extract directional cues from FOA channels, (ii) encode them into a latent representation, and (iii) model complex propagation phenomena including occlusion, reflections, and reverberation [liu2024audioldm].
The key challenge lies not only in localizing sound sources but also in handling multiple, dynamic sources within acoustically complex environments [schissler2016interactive]. By grounding audio generation in geometry- and material-aware acoustic descriptors, our model captures both static and dynamic elements of the scene. This enables real-time adaptation to source motion, ensuring accurate localization, separation, and a physically consistent audio field that reflects the spatial relationships inherent in 360-degree visual scenes [raghuvanshi2010precomputed].
@@ 2.3 Video Encoder
The Video Encoder extracts spatial and semantic cues from 360-degree video to support realistic sound propagation modeling and synchronized spatial audio rendering [tang2021learning]. It operates in three stages: (1) sound source localization and depth estimation, (2) semantic segmentation and scene reconstruction, and (3) feature extraction and multi-modal fusion.
@@ 2.3.1 Sound Source Localization and Depth Estimation
The encoder first detects and localizes sound-emitting objects in the scene [lin2024motion]. Each source $i$ is assigned a bounding box $\hat{b}_{i}$ and an activity score $\hat{y}_{i}$, optimized by:
@@
where $(u,v)$ are image coordinates, $D(u,v)$ is depth, and $\theta,\phi$ are the corresponding elevation and azimuth angles. The resulting point cloud serves as the geometric basis for acoustic modeling. To reconstruct the full scene, a hybrid approach combines Truncated Signed Distance Functions (TSDF) for large-scale structures with 3D Gaussian Splatting (3DGS) for fine details [curless1996volumetric, kerbl20233d], enabling accurate modeling of occlusion, reflections, and reverberation.
@@ 2.3.2 Semantic Segmentation and Scene Reconstruction
The encoder further applies semantic segmentation to classify scene elements (e.g., walls, floors, furniture). Each class is mapped to frequency-dependent acoustic material properties [van2014influence, ratnarajah2024listen2scene], enriching the reconstructed geometry with absorption and reflection parameters. These semantic and geometric cues are integrated into a 3D scene model [antani2012interactive], enabling simulation of occlusion, reflections, and reverberation based on both structure and material characteristics. This ensures that environmental effects such as sound blocking, scattering, and decay are faithfully captured.
@@ 2.3.3 Feature Extraction and Fusion
Finally, spatial and temporal features are extracted using Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) [courtney2019using]. These video-derived features are fused with audio representations to jointly model scene dynamics and acoustic conditions. The resulting multimodal features support real-time tracking of dynamic sound sources while incorporating propagation effects (occlusion, reflections, and reverberation), thereby enabling perceptually coherent and immersive spatial audio rendering.
@@ 2.4 Audio Encoder
The Audio Encoder processes FOA signals to extract directional, spectral, and temporal cues that align with the geometry- and material-aware representations produced by the Video Encoder. By coupling FOA features with reconstructed scene properties, it enables context-aware spatialization consistent with the physical layout and acoustic characteristics of the environment [zotter2019ambisonics, morgado2020learning].
@@ 2.4.1 FOA Extraction and Normalization
From the normalized channels, a CNN extracts compact representations of spectral and directional patterns. Convolutions over the time–frequency domain capture harmonic content and inter-channel correlations, while stacked layers aggregate these into higher-level spatial features [courtney2019using]. The resulting FOA embeddings form the basis for geometry-aware mapping and materially consistent spatial audio generation.
@@ 2.4.2 Spatial and Directional Mapping
To incorporate structural priors, we modulate FOA features with geometric distance and material-dependent absorption [zotter2019ambisonics]. This accounts for sound attenuation and redirection during propagation:
@@
Directional information is further captured by projecting FOA features onto a spherical harmonic basis, yielding a compact representation of spatial energy distributions [rafaely2015fundamentals]. This transformation reinforces alignment between audio embeddings and reconstructed scene geometry, enabling accurate reasoning about sound propagation across directions.
@@ 2.4.3 Saliency, Reverberation, and Output
To highlight perceptually relevant cues, an attention mechanism modulates FOA features with visual saliency and acoustic context [morgado2020learning]:
@@
Spatial realism is further enriched by augmenting FOA features with late reverberation profiles that capture long-range energy decay and diffusion [ratnarajah2024listen2scene]. Estimated from reconstructed geometry and material properties, these profiles complement direct sound and early reflections, yielding acoustically consistent reverberation patterns.
@@ 2.5 Conditional Diffusion Generator
Inspired from recent spatial audio diffusion frameworks such as Diff-SAGe [kushwaha2025diff] and ImmerseDiffusion [heydari2025immersediffusion], we employ a conditional diffusion model to synthesize FOA that remain consistent with reconstructed scene geometry and material properties. Operating in the latent FOA domain, the model integrates structural and acoustic cues—including occlusion, reflections, and reverberation—while accounting for dynamic and multiple sound sources. Starting from a noisy latent representation, a U-Net denoiser progressively reconstructs clean FOA signals over $T$ timesteps. Training follows the denoising diffusion probabilistic model (DDPM) objective:
@@ 2.5.1 Conditioning on Geometry and Material Properties
Reconstructed geometry and material attributes provide the foundation for physically grounded synthesis. The 3D mesh encodes structural layout and surface orientation, while material properties specify frequency-dependent absorption coefficients [ratnarajah2024listen2scene]. Embedding these features allows the model to account for attenuation, diffraction, and spatial filtering effects, thereby ensuring that the generated FOA is consistent with the reconstructed scene. Such conditioning is particularly important in scenarios involving dynamic or multiple sound sources, where accurate modeling of energy interaction with the environment is essential.
@@ 2.5.2 Conditioning for Occlusion, Reflections, and Reverberation
To capture realistic propagation effects, the model is further conditioned on occlusion, reflections, and reverberation. Occlusion features are derived from visibility analysis between listener and sources, modulated by material absorption. Early reflections are estimated by tracing geometric paths, providing echo-like cues that enhance spatial depth. Reverberation is represented using frequency-dependent $T_{60}(f)$ curves, which describe late decay characteristics [heydari2025immersediffusion]. Together, these cues enrich the conditioning stream, enabling FOA synthesis that incorporates both direct sound and its environmental response.